OpenStudy (anonymous):

HELP? The amount of interest earned in a savings account varies directly with the amount of money in the account. A certain bank offers a 2% savings rate. Write a direct variation equation for the amount of interest y earned on a balance of x. Then graph...

OpenStudy (anonymous):

Use the direct variance formula: y=kx.

OpenStudy (anonymous):

i need help on the graph part

OpenStudy (ranga):

y = kx. Here the interest rate is 2% = 2/100 = 0.02 So y = 0.02x Compare this to the equation of a straight line in the slope-intercept form: y = mx + b where m is the slope and b is the y-intercept. If you compare y = 0.02x and y = mx + b you can see the slope = 0.02 and the y-intercept b = 0. That means it is a straight line passing through the origin with a very small slope.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (ranga):